OpenStudy (anonymous):

Help please! Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = -5; y-intercept = (0, -6) slope = 5/3 ; y-intercept = (0, 2) slope = -3/5; y-intercept = (0, 2) slope =5/3 ; y-intercept = (0, -6) I picked the second one. Is that right?

OpenStudy (nurali):

-3y=-6-5x y=-6-5x/-3 y=5x/3+2 y=mx+b m=slope and b=y-intercept m=5/3;y-intercept (0,2)

OpenStudy (nurali):

if correct please click best response

OpenStudy (anonymous):

so my answer was right? The one I originally picked?

OpenStudy (nurali):

yes